A projectile has maximum range of 200m. What is the maximum height attained by it?


50 or 100m?

Grade:

1 Answers

Harish R
11 years ago

max range => angle =45 degrees

Range = u^2/g = 200.

Max height = 1/2*u^2 sin^2(45)/g = 200/2* 1/2 = 50 m.

Plz approve !!!!!!!!